Solve the following system of equations by graphing

x-2y=-4
x+2y=8

Answer by fractalier(6550)

Graphing is hard to do here, but I will give you some pointers...from
x-2y=-4
x+2y=8
put them into slope-intercept form...so
-2y = -x - 4
y = (1/2)x + 2
and
2y = -x + 8
y = (-1/2)x + 4
On the first one, plot a point at (0,2), then go up one and over two, and plot another point. Connect the dots.
On the second one, plot a point at (0,4), then go dwn one and over two, and plot another point. Connect the dots.
Where the lines intersect is the solution. Should be (2,3).